efren leaves home at 9 am and walks 4 miles per hour. his brother, Gregory leaves half an hour later and runs 8.5 miles per hour in the same direction. predict the time at which Gregory will catch up to efron.
Answer by richwmiller(17219)

Gaining speed 8.5 -4= 4.5 mph
Headstart 4*0.5 =2 miles
Time to catch up 2/4.5=0.44444444 hours
Time to catch up =26.6666667 minutes
Algebraic solution
4(t+0.5)=8.5t
4t+2.0=8.5t
2.0=8.5t-4t
2.0=4.5t
t=0.44444444 hours